# Bayesian Self-Distillation for Image Classification

### v1

@@ 4.2.4 Augmentation
CutMix. BSD consistently improves upon both the baseline and DLB across datasets and architectures. On CIFAR-100, BSD boosts convolutional networks by over 3 pp relative to the baseline, and by about +2.3+2.3 pp (DenseNet-169) and +1.3+1.3 pp (ResNet-50) compared to DLB. On TinyImageNet, BSD surpasses the baseline by roughly +2+2 pp for ResNet-101 and over +4+4 pp for DenseNet-201, both exceeding DLB by more than 1 pp. For ViT-B/16, BSD achieves gains of +0.5+0.5 pp over the baseline and +0.3+0.3 pp over DLB on CIFAR-100, and +1.15+1.15 pp and +0.6+0.6 pp on TinyImageNet. On CIFAR-10, improvements are smaller but consistent across all models. Compared with Table 1, CutMix enhances BSD’s performance, offering consistent gains over both standard training and DLB.
CutOut. The effect of CutOut is more nuanced. On CIFAR-10 and CIFAR-100, BSD consistently performs best, improving convolutional networks by more than +3+3 pp over the baseline and over +2+2 pp over DLB on CIFAR-100. For ViT-B/16, BSD remains competitive with the baseline and performs better than DLB, though the gains are modest given the already high baseline accuracy. On TinyImageNet, however, CutOut does not provide additional benefits: BSD still surpasses the baseline, but the accuracy is lower than without CutOut (cf. Table 1), and ViT-B/16 shows little change relative to baseline or DLB. We speculate that the poor results for BSD on TinyImageNet and DLB overall is due to over-regularization. Overall, BSD remains stronger than DLB under CutOut, but the augmentation itself appears less complementary to BSD than CutMix.
(a) Accuracy
(b) ECE
Figure 6: Impact of $\gamma$ and $c$. Validation accuracy and ECE for different values of $\gamma$ and $c$ (ResNet-50, CIFAR-100).
4.2.5 Ablation
@@
We perform ablations on the hyperparameters $\gamma$ and $c$, analyzing their effect on accuracy and ECE for ResNet-50 on CIFAR-100. The results are plotted in Figure 6. We observe high accuracy and low ECE for a range of values. Interestingly, the set of hyperparameters that minimizes ECE does not coincide exactly with the values that maximizes accuracy.
@@
The lowest ECE is achieved for $\gamma=0.97$ and $c=2000$, while the highest accuracy is observed at $\gamma=0.96$ and $c=1000$. Notably, the lowest ECE obtained is $1.33\%$ on the validation set, which is substantially lower than for the other methods in Table 2. While this comes at the cost of a slight decrease in accuracy, calibration may be prioritized for certain tasks or applications. Furthermore, by comparing $\gamma=1.0$ with the remainder of the values in Figure 6, we note that the discounting is a valuable addition that appears to improve both accuracy and ECE.
@@
While BSD appears offer low ECE and high accuracy for wide range of values of $\gamma$ and $c$, we note that if hyperparameters are chosen poorly (e.g. $\gamma=0.9$, $c=50.0$), performance deteriorates. This is likely due to underfitting caused by relying too much on predictions too early.
@@ 5 Discussion
Overall, our results demonstrate that BSD can improve test set accuracy, ECE and NLL across a variety of datasets, architectures, and augmentation strategies compared to conventional training and related self-distillation methods. BSD does not seem to overfit to noise in the same way as conventional training and contrastive self-distillation methods, and achieves higher test set accuracy under label noise than other architecture-preserving self-distillation methods. Additionally, BSD+ yields state-of-the-art accuracy under label noise on CIFAR-10.
@@ A.1 Experimental Details
(a) Conventional training
(b) BSD
Figure 7: Inter-class component $\bm{\mu}$ of dark knowledge. Semantical patterns emerge between classes, accentuated by BSD (ResNet-50, CIFAR-100).
(a) Conventional training
(b) BSD
Figure 8: Total sample-wise deviation $\delta(\mathbf{x}_{i})$ of dark knowledge. BSD promotes learning of sample-specific information (ResNet-50, CIFAR-100).
@@
Unless otherwise stated, all models trained on CIFAR and TinyImageNet datasets use the Adam optimizer (kingma2014adam) with a batch size of $256$, a maximum learning rate of $0.01$ scheduled via the 1cycle policy (smith2017superconvergence), and basic augmentations (random cropping and horizontal flipping) for 200 epochs (40 for ViTs, with a maximum learning rate of 5e-5). For ImageNet, we train with SGD for 90 epochs using a per-GPU batch size of $128$ on 8 GPUs, a learning rate of $0.5$ (scheduled via cosine annealing with 5 epochs of warm-up), a weight decay of 1e41e-4, and random resized crops of $224$ pixels combined with horizontal flipping.
@@
Architectures. Because of the small size of the images, the ResNet and DenseNet networks have been modified to include a $3\times 3$ convolution instead of the usual $7\times 7$ convolution for all datasets except ImageNet. The ViT-B (dosovitskiy2020image) model has been pretrained on ImageNet for all experiments and is available in Pytorch (paszke2019pytorch).
@@
Method-specific hyperparameters. For the temporal ensemble, we set the momentum parameter $\alpha=0.6$, gradually ramping up the distillation loss over the first 100 epochs (45 for ImageNet) and anneal Adam’s $\beta_{1}$ to zero during the final 50 epochs. For DLB, we follow shen2022self and use a temperature of $3$ and set $\alpha=1.0$, but train for only 100 epochs to offset the doubled mini-batch size. For PS-KD, we let $\alpha_{T}=0.8$ ($\alpha_{T}=0.3$ for ImageNet). For the knowledge distillation of the ensembles, we use a temperature of 3.
@@
Label Noise. Inspired by methods in semi-supervised learning, we construct BSD+ by combining BSD with a contrastive loss term. We do this by utilizing a strong and a weak set of augmentations, where the weak set is used for BSD, and the strong is used for the contrastive term. We define the contrastive loss as
$\mathcal{L}_{c}=\frac{\lambda_{a}}{m}\sum_{j=1}^{m}KL(f(T_{j}(\mathbf{x}_{i})),\hat{\mathbf{y}}^{t}),$ (11)
@@
where $m$ is the number of strongly augmented views, and $T_{j}$ is the corresponding transform. For the strong set of augmentations, we utilize AutoAugment (cubuk2018autoaugment), in combination with CutMix (yun2019cutmix) and Random Erasing (zhong2020random). For BSD+, we set $m=2$ for all noise levels and schedule the learning rate using cosine annealing (loshchilov2016sgdr).
@@
Because BSD may excessively smooth the label distributions when labels are noisy, we introduce a sharpening parameter $\tau$ into the loss
$\mathcal{L}\leftarrow\frac{1}{|B|k}\sum_{i\in B}\ell\left(\hat{\mathbf{y}}_{i}^{t},\frac{\left(\mathbf{y}_{i}^{t-1}\right)^{1/\tau}}{\|\left(\mathbf{y}_{i}^{t-1}\right)^{1/\tau}\|_{1}}\right),$ (12)
@@
for both BSD and BSD+, where $|B|$ is the batch size. In the absence of noise, we use $\tau=1$ to avoid promoting overconfident predictions, while we set $\tau=0.8$ for all experiments with label noise. For BSD, we set $\gamma=0.85$ and $c=2000$ when
(a) Conventional training
(b) BSD
Figure 9: Inter-class component $\bm{\mu}$ of dark knowledge. Semantical patterns emerge between classes, accentuated by BSD (ResNet-101, Tiny ImageNet).
(a) Conventional training
(b) BSD
Figure 10: Total sample-wise deviation $\delta(\mathbf{x}_{i})$ of dark knowledge. BSD promotes learning of sample-specific information (ResNet-101, Tiny ImageNet).
injecting symmetric noise, and $\gamma=0.9$ and $c=1000$ for asymmetric noise. For BSD+, we set $\gamma=0.95$ and $c=1000$ for all experiments, and set $\lambda_{c}=2$ under asymmetric noise. For BSD+ under symmetric noise we set $\lambda_{c}=4$, $\lambda_{c}=7$ and $\lambda_{c}=14$ for noise levels 20%, 50% and 80%, respectively.
@@ A.2 Additional Experiments
Emergence of dark knowledge. We include visualizations of the inter-class distributions for CIFAR-100 and Tiny ImageNet in Figures 7 and 9, with their corresponding sample-wise deviations plotted in Figures 8 and 10, respectively. We observe similar patterns as for ResNet-18 in Figure 2, but on a larger scale, and larger sample-wise deviations for the larger datasets.
To study the emergence of dark knowledge during training, we compute the average KL divergence between the output distributions of the model over epochs and those of the final model, while adjusting for temperature scaling (guo2017calibration). The results for are plotted in Figure 11, where the decrease of KL divergence over epochs suggest that dark knowledge is a property that emerges gradually.
Out-of-distribution calibration and detection. To evaluate model calibration under distributional shifts, we tested models trained on CIFAR-10 against the CIFAR-10-C benchmark, which applies 19 different corruptions (e.g., brightness, blur, noise) across five severity levels. As shown in Figure 12, BSD consistently achieves the lowest Expected Calibration Error (ECE) across all severity levels compared to the baseline and related methods. Importantly, the performance gap widens as data quality degrades, and we observe that the related methods’ ECE increases more rapidly at high corruption severities, while BSD maintains a flatter ECE curve. This indicates that BSD reduces overconfidence under increasing covariate shift.
Furthermore, we study the performance of BSD and related methods under domain shifts by measuring the area under the receiver operating characteristic curve (AUROC) for models trained on CIFAR-10 and CIFAR-100, evaluated against the Street View House Numbers (SVHN) dataset netzer2011reading. While the best performing method varies with dataset and model architecture, we observe that BSD yields the most consistent improvement across datasets and architectures. Notably, while TE exhibits instability under this shift on CIFAR-100 (e.g., degrading performance on DenseNet-169 trained on CIFAR-100 with respect to baseline), BSD maintains robust performance. For ViT-B, we observe performance saturation on CIFAR-10 (with all methods $98\%$), while for CIFAR-100, BSD yields a notable improvement ($2.29\%$) over the baseline.
@@
Figure 12: ECE under increasing corruptions. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD), on CIFAR-10 and evaluated on CIFAR-10-C.
Table 6: Out-of-Distribution Detection Performance (AUROC). Models trained on in-distribution datasets (CIFAR-10, CIFAR-100) are evaluated against SVHN as the out-of-distribution dataset. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD). The best results are highlighted in bold.
ID Dataset Model Baseline (%) TE (%) DLB (%) PS-KD (%) BSD (%)
@@
Table 7: Calibration Performance for WideResNet-40-1 on CIFAR-100. Result for all methods (excl. BSD and baseline) are from (hebbalaguppe2024calibration). The best results are highlighted in bold.
@@
BSD ($\gamma=0.97,c=2000$) $72.34{\scriptstyle\pm 0.26}$ $\mathbf{0.85}\,{\scriptscriptstyle\pm 0.08}$ $\mathbf{0.18}\,{\scriptscriptstyle\pm 0.00}$ $\mathbf{0.87}\,{\scriptscriptstyle\pm 0.19}$
BSD ($\gamma=0.96,c=3000$) $72.62\,{\scriptscriptstyle\pm 0.20}$ $3.28{\scriptstyle\pm 0.04}$ $0.19{\scriptstyle\pm 0.00}$ $3.19{\scriptstyle\pm 0.12}$
BSD ($\gamma=0.95,c=4000$) $\mathbf{72.79}\,{\scriptscriptstyle\pm 0.45}$ $7.08{\scriptstyle\pm 0.40}$ $0.23{\scriptstyle\pm 0.00}$ $7.08{\scriptstyle\pm 0.40}$
@@
(d) BSD
Figure 13: Penultimate layer representations of ResNet-50 on CIFAR-100. a-b: two semantically similar classes with one dissimilar class. c-d: Three semantically dissimilar classes.
Calibration. We train a WideResNet-40-1 on CIFAR-100 to benchmark BSD against different calibration methods including distillation-based method. We compare against Label Smoothing (LS) (szegedy2016rethinking), Temperature Scaling (TS) (guo2017calibration), MixUp (zhang2017mixup), Correctness Ranking Loss (CRL) (moon2020confidence), PS-KD (kim2021self), Multi-class Difference in Confidence and Accuracy (MDCA) (hebbalaguppe2022stitch), AdaFocal (ghosh2022adafocal), Calibration by Pairwise Constraints (CPC) (cheng2022calibrating), Margin-based Label Smoothing (MbLS) (liu2022devil), Adaptive and Conditional Label Smoothing (ACLS) (park2023acls) and combinations of the aforementioned method with knowledge distillation (hebbalaguppe2024calibration). The results in included in Table 7, where we observe that BSD yields the highest accuracy and the lowest ECE, Static Calibration Error (SCE) and Adaptive Calibration Error (ACE) of all methods.

### v2

@@ 4.1.4 Augmentation
ImageNet ResNet-152 $78.98{\scriptstyle\pm 0.14}$ $79.00{\scriptstyle\pm 0.02}$ $\mathbf{80.06}\,{\scriptscriptstyle\pm 0.10}$ $80.72{\scriptstyle\pm 0.18}$ $80.35{\scriptstyle\pm 0.03}$ $\mathbf{80.83}\,{\scriptscriptstyle\pm 0.07}$
DenseNet-264 $7.12{\scriptstyle\pm 0.05} 77$$7.75{\scriptstyle\pm 0.01}$ $\mathbf{78.91}\,{\scriptscriptstyle\pm 0.10}$ $79.02{\scriptstyle\pm 0.16}$ $79.46{\scriptstyle\pm 0.05}$ $\mathbf{79.57}\,{\scriptscriptstyle\pm 0.06}$
CutMix. DPS consistently improves upon both the baseline and DLB across datasets and architectures. The gains are most pronounced on CIFAR-100 and TinyImageNet, where the convolutional networks improve by between $1.7\text{--}4.1$ pp over the baseline and by between $1.1\text{--}2.3$ pp over DLB. On ImageNet, DPS beats the baseline and DLB by $0.1\text{--0.6}$ pp. For ViT-B/16, DPS achieves gains of +0.5+0.5 and +0.3+0.3 pp over the baseline DLB respectively on CIFAR-100, while improving by +1.15+1.15 and +0.6+0.6 pp on TinyImageNet.
CutOut. The effect of CutOut is more nuanced. On CIFAR-100, DPS consistently performs best, improving convolutional networks by more than +3+3 and +2+2 pp over the baseline and DLB, respectively. For ViT-B/16, DPS yields similar or slightly higher accuracy across datasets. While DPS surpasses the baseline on TinyImageNet, the accuracy is lower than without CutOut (cf. Table 1). We speculate CutOut is over-regularizing for the complex but low resolution dataset. On ImageNet, DPS increases accuracy on ImageNet over both the baseline and DLB by +1+1 and more than +2+2 pp, for ResNet-152 and DenseNet-264, respectively.
4.2 Analyzing Dark Knowledge
To analyze how DPS affects a model’s predicted distributions, we first validate that dark knowledge emerges naturally during training, and show that DPS promotes further discovery of underlying structures in the data. To formalize the notion of dark knowledge, we decompose the output of the network, for an input $\mathbf{x}_{i}$ with integer label $k=\underset{l}{\mathrm{argmax}}\ \mathbf{y}_{i,l}^{0}$, as
$f(\mathbf{x}_{i})=\bm{\mu}_{k}+\delta(\mathbf{x}_{i}).$ (9)
Here, $\bm{\mu}_{k}\in\Delta^{K}$ is the expected prediction of a given class
$\bm{\mu}_{k}=\mathbb{E}[f(X)\mid Y=k]$ (10)
and captures the inter-class component of dark knowledge. The function $\delta:\mathbb{R}^{d}\to\mathbb{R}^{K}$ captures the sample-specific deviation from $\bm{\mu}_{k}$. We approximate the expectation in Equation 10 using the empirical mean for a given class over the dataset, and compute $\delta(\mathbf{x}_{i})=f(\mathbf{x}_{i})-\bm{\mu}_{k}$. We hypothesize that the independent evidence counters resulting from the decomposition of the Dirichlet prior into independent gamma distributed random variables (Section 3.2) results in more nuanced predicted distributions for both classes and individual samples.
We visualize $\bm{\mu}$ and $\delta$ from a ResNet-18 trained on CIFAR-10 in Figures 5(a)-5(d). The log-scale heatmaps of $\bm{\mu}$ in Figures 5(a)-5(b) reveal that the networks learn meaningful inter-class relationships. For instance, they capture similarities between animals but also between bird and airplane, with these clusters being more prominent and the probabilities larger for DPS. The per-sample absolute deviation from $\bm{\mu}$ presented in Figures 5(c)-5(d) validates the emergence of $\delta$ and shows that DPS promotes learning of sample-specific information.
(a) Conventional training
(b) DPS
(c) Conventional training
(d) DPS
(e) Accuracy
(f) ECE
Figure 5: Visualization of dark knowledge and ablation study. (5(a), 5(b)) Inter-class component $\bm{\mu}$: Semantical patterns emerge, accentuated by DPS (ResNet-18, CIFAR-10). (5(c), 5(d)) Sample-specific component: DPS promotes learning of sample-specific information compared to conventional training. (5(e), 5(f)) Impact of $\gamma$ and $c$: Validation accuracy and ECE sensitivity (ResNet-50, CIFAR-100).
4.3 Ablation
@@
We perform ablations on the hyperparameters $\gamma$ and $c$, analyzing their effect on accuracy and ECE for ResNet-50 on CIFAR-100. The results are plotted in Figures 5(e) and 5(f). We observe high accuracy and low ECE for a range of values. Interestingly, the set of hyperparameters that minimizes ECE does not coincide exactly with the values that maximizes accuracy.
@@
The lowest ECE is achieved for $\gamma=0.97$ and $c=2000$, while the highest accuracy is observed at $\gamma=0.96$ and $c=1000$. Notably, the lowest ECE obtained is $1.33\%$ on the validation set, which is substantially lower than reported in Table 4.1.2. While this comes at the cost of a slight decrease in accuracy, calibration may be prioritized for certain tasks or applications. Furthermore, by comparing $\gamma=1.0$ with the remainder of the values in Figures 5(e) and 5(f), we note that the discounting is a valuable addition that appears to improve both accuracy and ECE.
@@
While DPS appears to offer low ECE and high accuracy for wide range of values of $\gamma$ and $c$, we note that if hyperparameters are chosen poorly (e.g., $\gamma=0.9$, $c=50.0$), performance deteriorates. This is likely due to underfitting caused by relying too much on predictions too early.
@@ 5 Discussion
Overall, our results demonstrate that DPS can improve test set accuracy, ECE and NLL across a variety of datasets, architectures, and augmentation strategies compared to conventional training and related self-distillation methods. DPS does not seem to overfit to noise in the same way as conventional training and contrastive self-distillation methods, and achieves higher test set accuracy under label noise than other architecture-preserving self-distillation methods. Additionally, DPS+ yields state-of-the-art accuracy under label noise on CIFAR-10.
@@ A.1 Experimental Details
Unless otherwise stated, all models trained on CIFAR and TinyImageNet datasets use the Adam optimizer (Kingma, 2014) with a batch size of $256$, a maximum learning rate of $0.01$ scheduled via the 1cycle policy (Smith and Topin, 2019), and basic augmentations (random cropping and horizontal flipping) for 200 epochs (40 for ViTs, with a maximum learning rate of 5e-5). For ImageNet, we train with SGD for $200$ epochs using a per-GPU batch size of $128$ on 8 GPUs, a learning rate of $0.5$ (scheduled via cosine annealing with 5 epochs of warm-up), a weight decay of 1e41e-4, and random resized crops of $224$ pixels combined with horizontal flipping.
@@
Architectures. Because of the small size of the images, the ResNet and DenseNet networks have been modified to include a $3\times 3$ convolution instead of the usual $7\times 7$ convolution for all datasets except ImageNet. The ViT-B (Dosovitskiy et al., 2020) model has been pretrained on ImageNet for all experiments and is available in Pytorch (Paszke et al., 2019).
@@
Method-specific hyperparameters. For the temporal ensemble, we set the momentum parameter $\alpha=0.6$, gradually ramping up the distillation loss over the first 100 epochs (45 for ImageNet) and anneal Adam’s $\beta_{1}$ to zero during the final 50 epochs (25 for ImageNet). For DLB, we follow Shen et al. (2022) and use a temperature of $3$ and set $\alpha=1.0$, but train for only 100 epochs to offset the doubled mini-batch size. For PS-KD, we let $\alpha_{T}=0.8$ ($\alpha_{T}=0.3$ for ImageNet). For the knowledge distillation of the ensembles, we use a temperature of 3.
@@
Label Noise. Inspired by methods in semi-supervised learning, we construct DPS+ by combining DPS with a contrastive loss term. We do this by utilizing a strong and a weak set of augmentations, where the weak set is used for DPS, and the strong is used for the contrastive term. We define the contrastive loss as
$\mathcal{L}_{c}=\frac{1}{m}\sum_{j=1}^{m}KL(f(T_{j}(\mathbf{x}_{i})),\hat{\mathbf{y}}^{t}),$ (11)
@@
where $m$ is the number of strongly augmented views, and $T_{j}$ is the corresponding transform, which is added to the supervised loss like $\mathcal{L}+\lambda_{c}\mathcal{L}_{c}$. For the strong set of augmentations, we utilize TrivialAugment (Müller and Hutter, 2021), in combination with CutMix (Yun et al., 2019) and Random Erasing (Zhong et al., 2020). For DPS+, we set $m=2$ for all noise levels and schedule the learning rate using cosine annealing (Loshchilov and Hutter, 2016).
@@
Because DPS may excessively smooth the label distributions when labels are noisy, we introduce a sharpening parameter $\tau$ into the loss
$\mathcal{L}\leftarrow\frac{1}{|B|}\sum_{i\in B}\ell\left(\hat{\mathbf{y}}_{i}^{t},\frac{\left(\mathbf{y}_{i}^{t-1}\right)^{1/\tau}}{\|\left(\mathbf{y}_{i}^{t-1}\right)^{1/\tau}\|_{1}}\right),$ (12)
@@
for both DPS and DPS+, where $|B|$ is the batch size. In the absence of noise, we use $\tau=1$ to avoid promoting overconfident predictions, while we set $\tau=0.8$ for all experiments with label noise. For DPS, we set $\gamma=0.85$ and $c=2000$ when injecting symmetric noise, and $\gamma=0.9$ and $c=1000$ for asymmetric noise. For DPS+, we set $\gamma=0.95$ and $c=1000$ for all experiments, and set $\lambda_{c}=2$ under asymmetric noise. For DPS+ under symmetric noise we set $\lambda_{c}=4$, $\lambda_{c}=7$ and $\lambda_{c}=14$ for noise levels 20%, 50% and 80%, respectively.
@@ A.2 Additional Experiments
Analyzing dark knowledge. We include visualizations of the inter-class distributions for CIFAR-100 and Tiny ImageNet in Figures 7 and 9, with their corresponding sample-wise deviations plotted in Figures 7 and 9, respectively. We observe similar patterns as for ResNet-18 in Figures 5(a)-5(b), but on a larger scale with increased sample-wise deviations.
(a) Conventional
(b) DPS
Figure 6: Inter-class component $\bm{\mu}$ of dark knowledge. Semantical patterns emerge between classes, accentuated by DPS (ResNet-50, CIFAR-100).
(c) Conventional
(d) DPS
Figure 7: Total sample-wise deviation $\delta(\mathbf{x}_{i})$. DPS promotes learning of sample-specific information (ResNet-50, CIFAR-100).
(a) Conventional
(b) DPS
Figure 8: Inter-class component $\bm{\mu}$ of dark knowledge. Semantical patterns emerge between classes, accentuated by DPS (ResNet-101, Tiny ImageNet).
(c) Conventional
(d) DPS
Figure 9: Total sample-wise deviation $\delta(\mathbf{x}_{i})$ of dark knowledge. DPS promotes learning of sample-specific information (ResNet-101, Tiny ImageNet).
@@
Figure 11: ECE under increasing corruptions. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS), on CIFAR-10 and evaluated on CIFAR-10-C.
Table 5: Out-of-Distribution Detection Performance (AUROC). Models trained on in-distribution datasets (CIFAR-10, CIFAR-100) are evaluated against SVHN as the out-of-distribution dataset. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS). The best results are highlighted in bold.
ID Dataset Model Baseline (%) TE (%) DLB (%) PS-KD (%) DPS (%)
@@
Table 6: Calibration Performance for WideResNet-40-1 on CIFAR-100. Result for all methods (excl. DPS and baseline) are from (Hebbalaguppe, 2024). The best results are highlighted in bold.
@@
DPS ($\gamma=0.97,c=2000$) $72.34{\scriptstyle\pm 0.26}$ $\mathbf{0.85}\,{\scriptscriptstyle\pm 0.08}$ $\mathbf{0.18}\,{\scriptscriptstyle\pm 0.00}$ $\mathbf{0.87}\,{\scriptscriptstyle\pm 0.19}$
DPS ($\gamma=0.96,c=3000$) $72.62\,{\scriptscriptstyle\pm 0.20}$ $3.28{\scriptstyle\pm 0.04}$ $0.19{\scriptstyle\pm 0.00}$ $3.19{\scriptstyle\pm 0.12}$
DPS ($\gamma=0.95,c=4000$) $\mathbf{72.79}\,{\scriptscriptstyle\pm 0.45}$ $7.08{\scriptstyle\pm 0.40}$ $0.23{\scriptstyle\pm 0.00}$ $7.08{\scriptstyle\pm 0.40}$
@@
(d) DPS
Figure 12: Penultimate layer representations of ResNet-50 on CIFAR-100. (a,b): Two semantically similar classes with one dissimilar class. (c,d): Three semantically dissimilar classes.
To study the emergence of dark knowledge during training, we compute the average KL divergence between the output distributions of the model over epochs and those of the final model, while adjusting for temperature scaling (Guo et al., 2017). The results are plotted in Figure 10, where the decrease of KL divergence over epochs suggest that dark knowledge is a property that emerges gradually.
Out-of-distribution calibration and detection. To evaluate model calibration under distributional shifts, we tested models trained on CIFAR-10 against the CIFAR-10-C benchmark, which applies 19 different corruptions (e.g., brightness, blur, noise) across five severity levels. As shown in Figure 11, DPS consistently achieves the lowest Expected Calibration Error (ECE) across all severity levels compared to the baseline and related methods. Importantly, the performance gap widens as data quality degrades, and we observe that the related methods’ ECE increases more rapidly at high corruption severities, while DPS maintains a flatter ECE curve. This indicates that DPS reduces overconfidence under increasing covariate shift.
Furthermore, we study the performance of DPS and related methods under domain shifts by measuring the area under the receiver operating characteristic curve (AUROC) for models trained on CIFAR-10 and CIFAR-100, evaluated against the Street View House Numbers (SVHN) dataset Netzer et al. (2011). While the best performing method varies with dataset and model architecture, we observe that DPS yields the most consistent improvement across datasets and architectures. Notably, while TE exhibits instability under this shift on CIFAR-100 (e.g., degrading performance on DenseNet-169 trained on CIFAR-100 with respect to baseline), DPS maintains robust performance. For ViT-B, we observe performance saturation on CIFAR-10 (with all methods $98\%$), while for CIFAR-100, DPS yields a notable improvement ($2.29\%$) over the baseline.
Calibration. We train a WideResNet-40-1 on CIFAR-100 to benchmark DPS against different calibration methods including distillation-based method. We compare against Label Smoothing (LS) (Szegedy et al., 2016), Temperature Scaling (TS) (Guo et al., 2017), MixUp (Zhang et al., 2017), Correctness Ranking Loss (CRL) (Moon et al., 2020), PS-KD (Kim et al., 2021), Multi-class Difference in Confidence and Accuracy (MDCA) (Hebbalaguppe et al., 2022), AdaFocal (Ghosh et al., 2022), Calibration by Pairwise Constraints (CPC) (Cheng and Vasconcelos, 2022), Margin-based Label Smoothing (MbLS) (Liu et al., 2022), Adaptive and Conditional Label Smoothing (ACLS) (Park et al., 2023) and combinations of the aforementioned method with knowledge distillation (Hebbalaguppe, 2024). The results are included in Table 6, where we observe that DPS yields the highest accuracy and the lowest ECE, Static Calibration Error (SCE) and Adaptive Calibration Error (ACE) of all methods.